OpenStudy (anonymous):

If you wanted a really steep hill for your roller coaster, would you use the graph of y = –2x2 + 6 or y = (-1/10)x2 + 6? Than create an equation that will be steeper than both equations and explain your reasoning.

OpenStudy (anonymous):

y=-2x^2+6 set X=10:: y=2*10^6+6 y=2*1000000 + 6 y = -1999994 y = (-1/10)x^2 + 6 set X=10:: y = (-1/10)10^2 + 6 y = (-10^2/10) + 6 y = (-10) + 6 y = -4

OpenStudy (anonymous):

Thus slope = -2 on the first one, slope is -1/10 on the second one.